OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30. intercept = (-2, 0); y-intercept = (0, 6) intercept = (0, -6); y-intercept = (2, 0) x-intercept = (2, 0); y-intercept = (0, -6) intercept = (0, 6); y-intercept = (-2, 0) My answer is x-intercept = (2, 0); y-intercept = (0, -6) Am I right? :)

OpenStudy (johnweldon1993):

That would be correct!

OpenStudy (anonymous):

You can set the equation to slope-intercept form. 15x-5y=30 -5y=-15x+30 y=3x-6 You are correct. 0=3x-6 6=3x x=2 y=3(0)-6 y=0-6 y=-6

OpenStudy (anonymous):

Thank so much! :)

OpenStudy (johnweldon1993):

No problem Good work!